# : Equal senior authors

### v1

@@ Abstract
Almost 30% of prostate cancer (PCa) patients undergoing radical prostatectomy (RP) experience biochemical recurrence (BCR), characterized by increased prostate specific antigen (PSA) and associated with increased mortality. Accurate early prediction of BCR, at the time of RP, would contribute to prompt adaptive clinical decision-making and improved patient outcomes. In this work, we propose prostate cancer BCR prediction via fused multi-modal embeddings (PROFUSEme), which learns cross-modal interactions of clinical, radiology, and pathology data, following an intermediate fusion configuration in combination with Cox Proportional Hazard regressors. Quantitative evaluation of our proposed approach reveals superior performance, when compared with late fusion configurations, yielding a mean C-index of 0.861 ($\sigma=0.112$) on the internal 5-fold nested cross-validation framework, and a C-index of 0.7103 on the hold out data of the CHIMERA 2025 challenge validation leaderboard.
@@ 3.2 Fusion Approaches
Fig. 1 (C) depicts the process of late fusion. To integrate the three modalities, we computed the median ($\overline{LP}$) of the predicted log-risk ($\overline{LP}_{P}$, $\overline{LP}_{R}$, $\overline{LP}_{C}$) from each of the independent CPH regressors to obtain the final patient-wise prediction. The late fusion is dependent on the different folds’ model training, in that we aggregated the models by taking the median (or mean) element-wise across the independent model weights from the 9 folds (the median or mean models step in Tab. 1), and used the aggregated model ($\overline{R}_{P}$, $\overline{R}_{R}$, $\overline{R}_{C}$) to predict the log-risk ($LR_{Pi}$, $LR_{Ri}$, $LR_{Ci}$, i=1,2,3,9i=1,2,3,...9) score from different features, where each pooled feature was attention-weighted and dependent on the original model before ensembling. The predicted TTR was calculated by $e^{(-\overline{LR})}$. For the internal data evaluation, we used $\times5$ nested cross-validation in Sec. 3.4, the aggregated models are calculated from the inner folds’ model weights since they share the same hold-out set.

### v2

@@ Abstract
Almost 30% of prostate cancer (PCa) patients undergoing radical prostatectomy (RP) experience biochemical recurrence (BCR), characterized by increased prostate specific antigen (PSA) and associated with increased mortality. Accurate early prediction of BCR, at the time of RP, would contribute to prompt adaptive clinical decision-making and improved patient outcomes. In this work, we propose prostate cancer BCR prediction via fused multi-modal embeddings (PROFUSEme), which learns cross-modal interactions of clinical, radiology, and pathology data, following an intermediate fusion configuration in combination with Cox Proportional Hazard regressors. Quantitative evaluation of our proposed approach reveals superior performance, when compared with late fusion configurations, yielding a mean C-index of 0.861 ($\sigma=0.112$) on the internal 5-fold nested cross-validation framework, and a C-index of 0.7107 on the hold out data of the CHIMERA 2025 challenge validation leaderboard.
@@ 3.2 Fusion Approaches
Fig. 1 (C) depicts the process of late fusion. To integrate the three modalities, we computed the median ($\overline{LR}$) of the predicted log-risk ($\overline{LR}_{P}$, $\overline{LR}_{R}$, $\overline{LR}_{C}$) from each of the independent CPH regressors to obtain the final patient-wise prediction. The late fusion is dependent on the different folds’ model training, in that we aggregated the models by taking the median (or mean) element-wise across the independent model weights from the 9 folds (the median or mean models step in Tab. 1), and used the aggregated model ($\overline{R}_{P}$, $\overline{R}_{R}$, $\overline{R}_{C}$) to predict the log-risk ($LR_{Pi}$, $LR_{Ri}$, $LR_{Ci}$, i=1,2,3,9i=1,2,3,...9) score from different features, where each pooled feature was attention-weighted and dependent on the original model before ensembling. The predicted TTR was calculated by $e^{(-\overline{LR})}$. For the internal data evaluation, we used $\times5$ nested cross-validation in Sec. 3.4, the aggregated models are calculated from the inner folds’ model weights since they share the same hold-out set.